A square garden has an area of 1024 squared feet, and the gardener wants to install a sprinkler (with a circluar spraying pattern) at the center of the garden. What is the minimum radius of spray the sprinkler would need in order to water all of the garden?
Answer by mananth(12270)

Area = 1024
side^2=1024
side = sqrt(1024)
side = 32
Find the diagonal
Half the doagonal is the radius to cover the entire square garden
32^2+32^2= diagonal ^2 ( Pythagoras theorem)
2048 = diagonal ^2
diagonal = 45.25
Radius = 45.25/2
radius = 22.625 feet